I am having trouble with figuring out how to solve this word problem. Please help me...
In a recent year, the estimated earnings for Jennifer Lopez was $2.5 million more than half the earnings for the band U2. If the total earnings were $106 million, what were the earnings for Jennifer Lopez and U2?

Answer by rfer(16322)

U2=x
J=0.5x+2.5 mill
---------------
x+0.5x+2.5=106
1.5x=106-2.5
1.5x=103.5
x=103.5/1.5
x=69 mill
0.5x+2.5 mill=37 mill